# Privacy-Preserving Logistic Regression Training with A Faster Gradient Variant

### v6

@@ 3.3 Quadratic Gradient Algorithms
This gradient variant, Quadratic Gradient, can be used to enhance various first-order gradient algorithms, such as NAG and Adagrad.
@@
$V_{t+1}$ $=\bm{\beta}_{t}+\alpha_{t}\cdot\nabla J(\bm{\beta}_{t}),$ (3)
$\bm{\beta}_{t+1}$ $=(1-\gamma_{t})\cdot V_{t+1}+\gamma_{t}\cdot V_{t},$ (4)
where $V_{t+1}$ is the intermediate variable used for updating the final weight $\bm{\beta}_{t+1}$ and $\gamma_{t}\in(0,1)$ is a smoothing parameter of moving average to evaluate the gradient at an approximate future position [11]. The enhanced NAG is to replace (3) with $V_{t+1}=\bm{\beta}_{t}+(1+\alpha_{t})\cdot G$. Our enhanced NAG method is described in Algorithm 1 .
Adagrad is a gradient-based algorithm suitable for dealing with sparse data. The updated operations of Adagrad and its quadratic-gradient version, for every parameter $\bm{\beta}_{[i]}$ at each iteration step $t$, are as follows, respectively:
$\bm{\beta}_{[i]}^{(t+1)}$ $=\bm{\beta}_{[i]}^{(t)}-\frac{\eta}{\epsilon+\sqrt{\sum_{k=1}^{t}\bm{g}_{[i]}^{(t)}\cdot\bm{g}_{[i]}^{(t)}}}\cdot\bm{g}_{[i]}^{(t)},$
$\bm{\beta}_{[i]}^{(t+1)}$ $=\bm{\beta}_{[i]}^{(t)}-\frac{1+\eta}{\epsilon+\sqrt{\sum_{k=1}^{t}G_{[i]}^{(t)}\cdot G_{[i]}^{(t)}}}\cdot G_{[i]}^{(t)}.$
@@
Figure 1: The training results of Adagrad and Enhanced Adagrad in the clear setting
Performance Evaluation We assess the performance of various algorithms in a clear setting using the Python programming language on a desktop computer with an Intel Core CPU G640 at 1.60 GHz and 7.3 GB RAM. Since our focus is on how fast the algorithms converge in the training phase, the loss function, maximum likelihood estimation (MLE), is selected as the only indicator. We evaluate four algorithms, NAG, Adagrad, and their quadratic-gradient versions (denoted as Enhanced NAG and Enhanced Adagrad, respectively) on the datasets that Kim et al. [11] adopted: the iDASH genomic dataset (iDASH), the Myocardial Infarction dataset from Edinburgh (Edinburgh), Low Birth weight Study (lbw), Nhanes III (nhanes3), Prostate Cancer study (pcs), and Umaru Impact Study datasets (uis). The genomic dataset is provided by the third task in the iDASH competition of 2017, which consists of 1579 records. Each record has 103 binary genotypes and a binary phenotype indicating if the patient has cancer. The other five datasets all have a single binary dependent variable. We also evaluate these algorithms on two large datasets from Han et al. [9]: a real financial data consisting of 422,108 samples over 200 features and the restructured public MNIST dataset that consists of 11,982 samples of the training dataset with 196 features. For a fair comparison with the baseline work [11], we use the learning rate $1+10/(1+t)$ since Kim et al. [11] select $10/(1+t)$ as their learning rate. In our experiments, we always use $\bar{H}$ = $-\frac{1}{4}X^{\intercal}X$ to construct our $\tilde{B}$ for the binary LR model.
Figures 1 and 2 show that except for the enhanced Adagrad method on the iDASH genomic dataset our enhanced methods all converge faster than their original ones in other cases. In all the Python experiments, the time to calculate the $\bar{B}$ in quadratic gradient $G$ before running the iterations and the time to run each iteration for various algorithms are not important and can be considered negligible. (few seconds).
@@
Figure 2: The training results of NAG and Enhanced NAG in the clear setting
@@
(a) The raw first-order gradient ascent algorithm with iterative formulas: $\bm{\beta}_{t+1}=\bm{\beta}_{t}+\mathrm{lr}\cdot\bm{g}$
(b) The raw quadratic gradient ascent algorithm with iterative formulas: $\bm{\beta}_{t+1}=\bm{\beta}_{t}+\mathrm{LR}\cdot G$
@@ 4 Homomorphic Training
Privacy-preserving logistic regression training based on HE techniques faces a difficult dilemma that no homomorphic schemes are capable of directly calculating the sigmoid function in the LR model. A common solution is to replace the sigmoid function with a polynomial approximation by using the widely adopted least-squares method. We can call a function named “ polyfit($\cdot$) ” in the Python package Numpy to fit the polynomial in a least-square sense. We adopt the degree 5 polynomial approximation $g(x)$ developed by Kim et al. [11], utilizing the least squares approach to approximate the sigmoid function over the interval $[-8,8]$: $g(x)=0.5+0.19131\cdot x-0.0045963\cdot x^{3}+0.0000412332\cdot x^{5}$
Given the training dataset $\text{X}\in\mathbb{R}^{n\times(1+d)}$ and training label $\text{Y}\in\mathbb{R}^{n\times 1}$, we adopt the same method that Kim et al. [11] used to encrypt the data matrix consisting of the training data combined with training-label information into a single ciphertext $\text{ct}_{Z}$. The weight vector $\beta^{(0)}$ consisting of zeros and the diagnoal elements of $\bar{B}$ are copied $n$ times to form two matrices. The data owner then encrypt the two matrices into two ciphertexts $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$, respectively. The ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ are as follows:
@@
where $\bar{B}_{[i][i]}$ is the diagonal element of $\bar{B}$ that is built from $-\frac{1}{4}X^{\intercal}X$.
The pulbic cloud takes the three ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ and evaluates the enhanced NAG algorithm to find a decent weight vector by updating the vector $\text{ct}_{\beta}^{(0)}$. Refer to [11] for a detailed description about how to calculate the gradient by HE programming.
@@
Edinburgh 1253 9 [11] 7 0.02 3.6 91.04 0.958
Ours 4 0.04 0.6 89.52 0.943
lbw 189 9 [11] 7 0.02 3.3 69.19 0.689
Ours 4 0.04 0.6 71.35 0.667
nhanes3 15649 15 [11] 7 0.16 7.3 79.22 0.717
Ours 4 0.31 4.5 79.23 0.637
pcs 379 9 [11] 7 0.02 3.5 68.27 0.740
Ours 4 0.04 0.6 63.20 0.733
uis 575 8 [11] 7 0.02 3.5 74.44 0.603
Ours 4 0.04 0.6 74.43 0.597
@@ 5 Experiments
For a fair comparison with [11], we utilized the same 10-fold cross-validation (CV) technique on the same iDASH dataset consisting of 1579 samples with 18 features and the same 5-fold CV technique on the other five datasets. Like [11], We consider the average accuracy and the Area Under the Curve (AUC) as the main indicators. Tables 1 and 2 display the results of the two experiments, respectively. The two tables also provide the average evaluation running time for each iteration and the storage (encrypted dataset for the baseline work and encrypted dataset and $\bar{B}$ for our method). We adopt the same packing method that Kim et al. [11] proposed and hence our solution has similar storage of ciphertexts to [11] with some extra ciphertexts to encrypt the $\bar{B}$. We chose $1+0.9^{t}$ as our learning rate configuration.
The parameters of HEAAN we set are same to [11]: $logN=16$, $logQ=1200$, $logp=30$, $slots=32768$, which ensure the security level $\lambda=80$. We use a larger $logp=40$ to encrypt the matrix ${\bar{B}}$ for preserving the precision of ${\bar{B}}$. Refer to [11] for details on these parameters. Since our enhanced NAG method need one more ciphertext multiplication than the baseline work, consuming more modulus, our solution thus can only perform $4$ iterations of the enhanced NAG method. Yet despite only $4$ iterations, our enhanced NAG method still produces a comparable result.

### v7

@@ 3.3 Quadratic Gradient Algorithms
This gradient variant, Quadratic Gradient, can be used to enhance various first-order gradient algorithms:
  1. 1.
@@
    $V_{t+1}$ $=\boldsymbol{\beta}_{t}+\eta_{t}\cdot\nabla J(\boldsymbol{\beta}_{t}),$ (3)
    $\boldsymbol{\beta}_{t+1}$ $=(1-\gamma_{t})\cdot V_{t+1}+\gamma_{t}\cdot V_{t},$ (4)
    where $V_{t+1}$ is the intermediate variable used for updating the final weight $\boldsymbol{\beta}_{t+1}$ and $\gamma_{t}\in(0,1)$ is a smoothing parameter of moving average to evaluate the gradient at an approximate future position (Kim et al., 2018a). The enhanced NAG is to replace (3) with $V_{t+1}=\boldsymbol{\beta}_{t}+N_{t}\cdot G,$ where $N_{t}$ is the new learning rate for the enhanced algorithm usually setted to $1+\eta_{t}$. Our enhanced NAG is described in Algorithm 1 .
  2. 2.
    Adagrad is a gradient-based algorithm suitable for dealing with sparse data. The updated operations of Adagrad and its quadratic-gradient version, for every parameter $\boldsymbol{\beta}_{[i]}$ at each iteration step $t$, are as follows, respectively:
    $\boldsymbol{\beta}_{[i]}^{(t+1)}$ $=\boldsymbol{\beta}_{[i]}^{(t)}-\frac{\eta_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}\boldsymbol{g}_{[i]}^{(t)}\cdot\boldsymbol{g}_{[i]}^{(t)}}}\cdot\boldsymbol{g}_{[i]}^{(t)},$
    $\boldsymbol{\beta}_{[i]}^{(t+1)}$ $=\boldsymbol{\beta}_{[i]}^{(t)}-\frac{N_{t}}{\epsilon+\sqrt{\sum_{k=1}^{t}G_{[i]}^{(t)}\cdot G_{[i]}^{(t)}}}\cdot G_{[i]}^{(t)}.$
  3. 3.
    Adam is an optimization algorithm that combines the benefits of Adagrad, which adapts learning rates for each parameter, with momentum, allowing it to converge faster and handle sparse gradients more effectively. Similar to how Adagrad transitions to its enhanced versions, the update formula for enhanced Adam, also employs a new learning rate and replaces the gradient with its quadratic form.
@@
Figure 1: The training results of NAG and Enhanced NAG in the clear
@@
Figure 2: The training results of Adagrad and Enhanced Adagrad in the clear
(a) The iDASH dataset
(b) The Edinburgh dataset
(c) The lbw dataset
(d) The nhanes3 dataset
(e) The pcs dataset
(f) The uis dataset
(g) restructured MNIST dataset
(h) The private financial dataset
Figure 3: The training results of Adam and Enhanced Adam in the clear
Performance Evaluation We evaluate the performance of various algorithms in a non-encrypted context using Python on a desktop computer with an Intel Core G640 CPU at 1.60 GHz and 7.3 GB of RAM. Given that our focus is on the convergence speed of the algorithms during training, we use the loss function, specifically maximum likelihood estimation (MLE), as the only indicator. We evaluate six algorithms—NAG, Adagrad, Adam, and their quadratic-gradient variants (denoted as Enhanced NAG, Enhanced Adagrad, and Enhanced Adam, respectively)—using the datasets adopted by Kim et al. (2018a): the iDASH genomic dataset (iDASH), the Myocardial Infarction dataset from Edinburgh (Edinburgh), the Low Birth Weight Study (lbw), Nhanes III (nhanes3), the Prostate Cancer Study (pcs), and the Umaru Impact Study datasets (uis). The genomic dataset from the third task of the 2017 iDASH competition comprises 1,579 records, each featuring 103 binary genotypes and a binary phenotype indicating whether the patient has cancer. The remaining five datasets each contain a single binary dependent variable. We also evaluate these algorithms on two large datasets from Han et al. (2019): a real financial dataset with 422,108 samples and 200 features, and the restructured public MNIST dataset, which includes 11,982 samples from the training set with 196 features. For a fair comparison with the baseline work (Kim et al., 2018a), we use the learning rate $1+\frac{10}{1+t}$, in line with Kim et al.’s choice of $\frac{10}{1+t}$ for their learning rate. In our experiments, we consistently use $\bar{H}=-\frac{1}{4}X^{\top}X$ to construct $\tilde{B}$ for the binary logistic regression model.
Figures 21, and 3 show that, except for the Enhanced Adagrad and Adam methods on the iDASH genomic dataset, our enhanced methods converge faster than their original counterparts in all other cases. All figures demonstrate that the enhanced algorithms converge empirically. In all the Python experiments, the time required to calculate $\bar{B}$ in the quadratic gradient $G$ before running the iterations, as well as the time to execute each iteration for various algorithms, is minimal and can be considered negligible (a few seconds).
@@
(a) The raw first-order gradient ascent algorithm with iterative formulas: $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\mathrm{lr}\cdot\boldsymbol{g}$
(b) The raw quadratic gradient ascent algorithm with iterative formulas: $\boldsymbol{\beta}_{t+1}=\boldsymbol{\beta}_{t}+\mathrm{LR}\cdot G$
@@ 4 Homomorphic Training
Privacy-preserving logistic regression training based on HE techniques faces a difficult dilemma that no homomorphic schemes are capable of directly calculating the sigmoid function in the LR model. A common solution is to replace the sigmoid function with a polynomial approximation by using the widely adopted least-squares method. We can call a function named “ polyfit($\cdot$) ” in the Python package Numpy to fit the polynomial in a least-square sense. We adopt the degree 5 polynomial approximation $g(x)$ developed by Kim et al. (2018a), utilizing the least squares approach to approximate the sigmoid function over the interval $[-8,8]$: $g(x)=0.5+0.19131\cdot x-0.0045963\cdot x^{3}+0.0000412332\cdot x^{5}$
Given the training dataset $\text{X}\in\mathbb{R}^{n\times(1+d)}$ and training label $\text{Y}\in\mathbb{R}^{n\times 1}$, we adopt the same method that Kim et al. (2018a) used to encrypt the data matrix consisting of the training data combined with training-label information into a single ciphertext $\text{ct}_{Z}$. The weight vector $\beta^{(0)}$ consisting of zeros and the diagnoal elements of $\bar{B}$ are copied $n$ times to form two matrices. The data owner then encrypt the two matrices into two ciphertexts $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$, respectively. The ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ are as follows:
@@
where $\bar{B}_{[i][i]}$ is the diagonal element of $\bar{B}$ that is built from $-\frac{1}{4}X^{\top}X$.
The pulbic cloud takes the three ciphertexts $\text{ct}_{Z}$, $\text{ct}_{\beta}^{(0)}$ and $\text{ct}_{\bar{B}}$ and evaluates the enhanced NAG algorithm to find a decent weight vector by updating the vector $\text{ct}_{\beta}^{(0)}$. Refer to (Kim et al., 2018a) for a detailed description about how to calculate the gradient by HE programming.
@@
Edinburgh 1253 9 Ours 5 4 0.04 0.6 89.52 0.943
(Kim et al., 2018a) 5 7 0.02 3.6 91.04 0.958
lbw 189 9 Ours 5 4 0.04 0.6 71.35 0.667
(Kim et al., 2018a) 5 7 0.02 3.3 69.19 0.689
nhanes3 15649 15 Ours 5 4 0.31 4.5 79.23 0.637
(Kim et al., 2018a) 5 7 0.16 7.3 79.22 0.717
pcs 379 9 Ours 5 4 0.04 0.6 63.20 0.733
(Kim et al., 2018a) 5 7 0.02 3.5 68.27 0.740
uis 575 8 Ours 5 4 0.04 0.6 74.43 0.597
(Kim et al., 2018a) 5 7 0.02 3.5 74.44 0.603
@@ 5 Experiments
For a fair comparison with (Kim et al., 2018a), we utilized the same 10-fold cross-validation (CV) technique on the same iDASH dataset consisting of 1579 samples with 18 features and the same 5-fold CV technique on the other five datasets. Like (Kim et al., 2018a), We consider the average accuracy and the Area Under the Curve (AUC) as the main indicators. Tables 1 and 2 display the results of the two experiments, respectively. The two tables also provide the average evaluation running time for each iteration and the storage (encrypted dataset for the baseline work and encrypted dataset and $\bar{B}$ for our method). We adopt the same packing method that Kim et al. (2018a) proposed and hence our solution has similar storage of ciphertexts to (Kim et al., 2018a) with some extra ciphertexts to encrypt the $\bar{B}$. We chose $1+0.9^{t}$ as our learning rate configuration.
The parameters of HEAAN we set are same to (Kim et al., 2018a): $logN=16$, $logQ=1200$, $logp=30$, $slots=32768$, which ensure the security level $\lambda=80$. We use a larger $logp=40$ to encrypt the matrix ${\bar{B}}$ for preserving the precision of ${\bar{B}}$. Refer to (Kim et al., 2018a) for details on these parameters. Since our enhanced NAG method need one more ciphertext multiplication than the baseline work, consuming more modulus, our solution thus can only perform $4$ iterations of the enhanced NAG method. Yet despite only $4$ iterations, our enhanced NAG method still produces a comparable result.
@@ 6 Conclusion
Impact Statements
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.